# Scalar-Induced Gravitational Waves from self-resonant preheating in $\bm{\alpha}$-attractor models

### v1

@@ Abstract
After the inflationary phase, the universe enters the preheating phase, during which the inflaton field rolls down its potential and oscillates. When the potential significantly deviates from a parabolic shape at its minimum, these oscillations trigger an instability in the scalar perturbations, leading to their amplification. This phenomenon, known as self-resonance, has important implications in cosmology. Notably, since scalar perturbations couple to tensor perturbations at second order in the equations of motion, this amplification results in the production of Gravitational Waves (GWs), referred to as Scalar-Induced Gravitational Waves (SIGWs). In this study, we investigate the production of SIGWs during the preheating phase for a class of inflationary models known as $\alpha$-attractors, characterized by a single parameter $\alpha$. We focus on small values of this parameter, specifically $\alpha\sim\mathcal{O}(10^{-1}-10^{-4})$, where the self-resonance effect is particularly pronounced. We obtain lower bounds on this parameter, $\log_{10}(\alpha)>-3.54$ for the T-model and $\log_{10}(\alpha)>-3.17$ for the E-model, based on the energy density of SIGWs constrained by Big Bang nucleosynthesis, which ultimately translates into lower bounds on the tensor-to-scalar ratio, $r>9.61\times 10^{-7}$ for the T-model and $r>2.25\times 10^{-6}$ for the E-model.
1 Introduction
After the Planck 2013-2018 data, there has been a surge in the building of single-field Starobinsky-like inflationary scenarios Linde:2014nna in various frameworks of quantum gravity due to stringent constraints on multifield inflation. The vanilla models of single-field inflation predict the scalar spectral index ($n_{s}$) and tensor-to-scalar ratio ($r$) as Kehagias:2013mya ; Ellis:2013nxa
@@ 1 Introduction
where $\alpha$ is a parameter that depends on the framework for the ultra-violet (UV) completion of gravity. In the context of $\alpha$-attractor inflation, the $\alpha$ parameter indicates the curvature of the Kähler geometry in supergravity Kallosh:2013yoa ; Kallosh:2013daa ; Kallosh:2013hoa ; Kaiser:2013sna ; Kallosh:2015lwa ; Iacconi:2023mnw ; Carrasco:2015uma ; Kallosh:2015zsa . It is important to note that models of $\alpha$-attractors could predict, in general, any smaller value for the tensor-to-scalar ratio compatible with the latest bound $r<0.032$ BICEP:2021xfz ; Tristram:2021tvh . Although future observations aim to detect primordial gravitational waves, there is a lack of a theoretical lower bound on the value of $\alpha$ except through considerations of reheating Iacconi:2023mnw . However, right after inflation ends, the universe enters a phase known as preheating Kofman:1994rk ; Kofman:1997yn during which the universe behaves effectively as matter dominated Jerome:2020 . Furthermore, when the inflationary potential deviates significantly from quadratic behavior (as is the case for small $\alpha$), the oscillations trigger the phenomenon of self-resonance del-Corral:2024vcm ; Hertzberg:2014iza ; Hertzberg:2014jza ; Jedamzik:2010dq ; Martin:2019nuw ; Sfakianakis:2018lzf ; Ballesteros:2024hhq ; Martin:2020fgl ; Shafi:2024jig ; Mahbub:2023faw ; Sang:2020kpd ; Zhang:2023hjk ; Child:2013ria , which translates into a fast and substantial amplification of scalar perturbations that act as a source of Gravitational Waves (GWs), since these fluctuations couple to tensor ones at second order. The GWs produced in this fashion are called Scalar-Induced Gravitational Waves (SIGWs) Guzzetti:2016 ; Bauman:2009 ; Maggiore:2007 ; Dodelson:2003 ; Domenech:2021 ; Gravitation:1973 ; Franciolini:2021 ; Baumann:2007 ; Ananda:2006 ; Assadullahi:2009 , and can be so efficiently produced to reach the Big Bang Nucleosynthesis (BBN) bound. This precise fact is what our investigation utilizes to derive a new lower bound on $\alpha$, based solely on the considerations of preheating instabilities and the consequent generation of SIGWs. Our result is robust even for a small duration of preheating of 1-2 e-folds, as one can realistically expect after the end of inflation.
@@
Recently, we have shown that these instabilities during reheating could lead to the generation of primordial black holes (PBH) del-Corral:2023apl ; draft-alpha-PBH in Starobinsky as well as in $\alpha$-attractor inflationary models. However, the study in this paper focuses on SIGWs produced irrespective of whether or not PBH are generated during preheating. It is also important to note that SIGWs are a secondary effect. At the linearized level, both scalar and tensor perturbations decouple. Therefore, we compute the linearized effect, which we call background GWs (BGWs), and the non-linear (secondary) effect, i.e., the SIGWs.
@@
Figure 1: Schematic representation of the production of SIGWs during self-resonant preheating. The upper panel shows the evolution of selected comoving scales as a function of the scale factor of the universe, as well as the comoving Hubble radius $(aH)^{-1}$. The suffixes CMB, end, and rh correspond to the CMB pivot scale, the end of inflation, and the beginning of reheating, respectively. The evolution of $\Phi_{k}$ in the bottom panel is shown schematically to illustrate the approximate behavior of the perturbations. See the text for details.
A schematic representation of the scenario described above is shown in Fig. 1. The key aspect is the definition of the relevant scales involved in the instability, labeled as band of instability. This band is bounded by two characteristic scales: $k_{L}$ (lower) and $k_{U}$ (upper). The lower bound, $k_{L}$, corresponds to the last mode entering the horizon during preheating and is defined as $k_{L}=a(t_{\text{rh}})H(t_{\text{rh}})$, where $a(t)$ is the scale factor of the universe, $H(t)$ the Hubble rate, and $t_{\text{rh}}$ marks the end of the preheating phase. The upper bound, $k_{U}$, denotes the largest mode affected by self-resonance and is determined using Floquet theory del-Corral:2024vcm . For a sufficiently long preheating, the quadratic approximation is good enough so that $k_{U}$ can be defined as follows Jerome:2020
$\left(\frac{k_{U}}{a}\right)^{-1}=\frac{1}{M}\left(\frac{2}{3\phi_{\text{end}}}\left(\frac{a}{a_{\text{end}}}\right)^{3}\right)^{1/4}.$ (2)
@@
Here $M$ is the mass of the scalar field $\phi$, and the suffix “${}_{\text{end}}$” stands for evaluation at the end of inflation. When the modes enter the self-resonance instability (green shaded area), the scalar perturbations, such as $\Phi_{\bm{k}}$, amplify. Their behavior is illustrated in the bottom panel of Fig. 1 for two distinct scales, $k_{1}$ and $k_{2}$, separated by the scale $k_{\text{end}}=a_{\text{end}}H_{\text{end}}$, defined as the last scale to exit the horizon during inflation. The first, shown in red, remains approximately constant after re-entering the Hubble horizon $(aH)^{-1}$ during preheating. Although this may appear as non-amplification, it actually reflects the counteraction of the usual decay of the gravitational potential at sub-horizon scales in an expanding universe, ensuring $\Phi_{\bm{k}}$ remains stable during this phase. The mode $k_{2}$, shown in blue, experiences stronger amplification as it spends more time within the region of instability of self-resonance. It should be noted that the Floquet exponent, responsible for amplification, is not the same for all modes within the band of instability del-Corral:2024vcm . In fact, the amplification is particularly stronger for modes $k>k_{\text{end}}$. During this amplification process, SIGWs are generated, and after that, the SIGWs freely propagate throughout the universe with minimal interaction with matter.
@@
Depending on the specific range of scales affected by the self-resonance, the SIGWs will have an imprint in a particular span of frequencies. In general, these scales are small and around the last scale to exit the horizon during inflation, which implies that we lie on the Very-High-Frequency (VHF) region of the GW spectrum (100 kHz – 1 THz) Kuroda:2015owv ; Aggarwal:2020olq ; Page:2020zbr ; Dandoy:2024oqg . This VHF region has been explored in the literature in the contexts of GUT scale PBH and quantum gravity Anantua:2008am ; Zagorac:2019ekv ; Addazi:2024qcv . Although some detectors (in operation and/or planned) work on the VHF band of the GW spectrum, their sensitivities are too low to detect a stochastic background of SIGWs. These sensitivities are shown in Fig. 2 in terms of the fractional energy density of gravitational waves, $\Omega_{\text{GW}}$. One can observe that all of them lie above the cosmological bound coming from Big Bang nucleosynthesis (BBN) Smith:2006nka ; Maggiore:1999vm , making it impossible to detect a stochastic background of GWs formed during preheating (otherwise, it would conflict with BBN predictions). In the VHF band, one of interest in this work, the most promising probes to detect GWs are:
  • Optically levitated dielectric sensors: These are based on the use of optically trapped and cooled dielectric microspheres or microdisks that experience a force when a GW passes through. A 1-meter prototype, the Optically levitated sensor (OLS) Aggarwal:2020umq , is under construction at Western University in the USA.
  • Bulk acoustic wave (BAW) devices: The vibrations of a resonant mass detector due to the passage of a GW could be read through the piezoelectric effect and Superconducting Quantum Interference Devices (SQUIDs) Goryachev:2014yra . A BAW device has been running since 2018 at the University of Western Australia Goryachev:2021zzn ; Goryachev:2014yra .
  • Interferometers up to 100 MHz: The quantum Cramér-Rao bound implies that, for a given laser power, higher bandwidth is needed to increase the sensitivity Mizuno:1995iqa . This implies that a broadband interferometer with the VIRGO, LIGO, or KAGRA-level sensitivity is not viable in the MHz region. However, there are several other interferometers in this bandwidth, such as the Holometer experiment at Fermilab holometer . It consists of two co-located power-recycled Michelson interferometers with 40-meter-long arms, which increase sensitivity in the search for stochastic GWs.
  • Gertsenshtein effect: This consists of the conversion of photons into GWs (and vice versa) in the presence of a magnetic field. Although these kinds of detectors are not yet built, Ref. Aggarwal:2020olq points out the possibility of using existing experiments originally designed to search for axion-like particles from the Sun, such as IAXO Armengaud:2014gea . Ref. Ringwald:2020ist explores this possibility using Single Photon Detectors (SPD) and Heterodyne radio receivers (HET). Another possibility is to rely on observations of large-scale cosmic regions, where magnetic fields may extend over kpc to Mpc scales, thereby increasing the effective detector volume. Thus, using the astrophysical data from radio telescopes, such as EDGES or ARCADE Domcke:2020yzq ; Lella:2024dus , one can constrain $\Omega_{\text{GW}}$.
Apart from the VHF band, we also show for comparison in Fig. 2 the sensitivity curves of GW detectors in lower frequency bands: the Very-Low-Frequency (VLF), the Low-Frequency (LF), the Middle-Frequency (MF), and the High-Frequency (HF). See Kuroda:2015owv ; Aggarwal:2020umq and references therein for details about the detectors and the acronyms. The two curves labeled as $r=0.028$ and $r=2.25\times 10^{-6}$ in Fig. 2 correspond to the energy density of GWs (background $+$ induced) from an E-model with $\alpha=8.4$ and $\alpha=10^{-3.17}$, respectively, where $r$ is the tensor to scalar ratio (1). The upper bound on $r$ corresponds to the one obtained using 10 datasets from the BICEP/Keck Array 2015 and 2018, Planck releases 3 and 4, and LIGO-Virgo-KAGRA Collaboration Galloni:2022mok . The lower bound is obtained in this work using SIGWs and the BBN bound. See Sec. 3.2 for details.
@@
Figure 2: Sensitivity curves of some GW detectors (in terms of $\Omega_{GW}$) as a function of the frequency in Hertz. Also shown is the BBN bound, taken from Smith:2006nka ; Maggiore:1999vm . Solid (dashed) lines represent operative (planned or theorized) GW detectors. The figure is produced from the analysis of Kuroda:2015owv ; Aggarwal:2020olq and references therein. The bottom lines with arrows mark the limits of the standard frequency bands of GWs, as defined in Kuroda:2015owv . The blue and magenta curves labeled as $r=0.028$ and $r=2.25\times 10^{-6}$ represent two spectra of GWs computed from the upper and lower bounds on the tensor-to-scalar ratio; see the text for details. These are computed using the method outlined in this work.
@@
The inflationary models we study in this work are the well-known $\alpha$-attractor models Kallosh:2013yoa ; Kallosh:2013daa ; Kallosh:2013hoa ; Kaiser:2013sna ; Kallosh:2015lwa ; Iacconi:2023mnw ; Carrasco:2015uma ; Kallosh:2015zsa , which exhibit excellent agreement with Planck data Planck:2018jri and have universal predictions of observables. For this reason, they have gained a lot of attention. There are two main subclasses of $\alpha$-attractor models, called T-models and E-models, named by the shape of the inflaton potentials (see Fig. 3), which are given by
@@
$V_{T}(\phi)=3\alpha M^{2}M_{\mathrm{Pl}}^{2}\tanh^{2}\left[\frac{1}{\sqrt{6\alpha}}\phi\right],\quad V_{E}(\phi)=\frac{3\alpha M^{2}M_{\mathrm{Pl}}^{2}}{4}\left(1-e^{-\sqrt{\frac{2}{3\alpha}}\phi}\right)^{2}$ (3)
@@
where $M_{\mathrm{Pl}}$ is the Planck mass and $M=1.3\times 10^{-5}M_{\mathrm{Pl}}$ is fixed by the amplitude of primordial power spectra. Beyond their inflationary observables (1), we show that these models are also particularly interesting from the perspective of SIGWs. As the value of $\alpha$ decreases, the minimum of the potential (at $\phi=0$) starts to deviate from the quadratic (parabola) behavior, triggering the self-resonance phenomenon and amplifying the scalar perturbations. Fig. 3 compares both potentials for several values of $\alpha$, revealing an evident asymmetry in the E-model case. This, as shown in del-Corral:2024vcm , makes the self-resonance effect stronger in this class of potentials, ultimately imposing a tighter constraint on $\alpha$ for the E-models and, consequently, on the tensor-to-scalar ratio. These differences will be further analyzed in Sec. 3, where their impact on tensor perturbations is clarified.
@@
This work is organized as follows: In Sec. 2, we analyze the evolution of scalar metric perturbations in a perturbed flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric. In Sec. 3, we examine the generation of SIGWs from these scalar perturbations and numerically compute the energy density of GWs for different values of $\alpha$. Finally, conclusions are drawn in Sec . 4, and Appendixes A and B show some complementary computations, such as the derivation of the equation of motion for the SIGWs or its power spectrum. Throughout the paper, we follow the metric signature ($+$)(-$+$)( - + + + ) and $\hbar=c=1$ with reduced Planck mass $M_{\rm Pl}^{2}=\frac{1}{8\pi G}$.
2 Scalar perturbations
@@
Figure 3: Normalized $\alpha$-attractor potentials for (a) T-model and (b) E-model plotted for several values of $\alpha$: $10^{-1}$ in blue, $10^{-2}$ in orange, $10^{-3}$ in green and $10^{-4}$ in red.

### v2

@@ Abstract
After the inflationary phase, the universe enters the preheating phase, during which the inflaton field rolls down its potential and oscillates. When the potential significantly deviates from a parabolic shape at its minimum, these oscillations trigger an instability in the scalar perturbations, leading to their amplification. This phenomenon, known as self-resonance, has important implications in cosmology. Notably, since scalar perturbations couple to tensor perturbations at second order in the equations of motion, this amplification results in the production of Gravitational Waves (GWs), referred to as Scalar-Induced Gravitational Waves (SIGWs). In this study, we investigate the production of SIGWs during the preheating phase for a class of inflationary models known as $\alpha$-attractors, characterized by a single parameter $\alpha$. We focus on small values of this parameter, specifically $\alpha\sim\mathcal{O}(10^{-1}-10^{-4})$, where the self-resonance effect is particularly pronounced. We obtain lower bounds on this parameter, $\log_{10}(\alpha)>-3.71(-3.51)$ for the T-model and $\log_{10}(\alpha)>-3.48(-3.19)$ for the E-model, and assuming a duration of preheating of 1 e-folds (5 e-folds). These bounds are obtained based on the energy density of SIGWs constrained by Big Bang nucleosynthesis, which ultimately translates into lower bounds on the tensor-to-scalar ratio, $r>6.54\times 10^{-7}(1.03\times 10^{-6})$ for the T-model and $r>1.11\times 10^{-6}(2.14\times 10^{-6})$ for the E-model. Note that these bounds on $\alpha$ and $r$ are derived within the linear framework of tensor fluctuations at the level of equations of motion, which nevertheless include scalar-scalar-tensor interactions with metric and matter fields. However, fully non-linear approaches, with all higher-order metric fluctuations, would be needed in the future to further validate these conclusions.
preprint: APS/123-QED
I Introduction
After the Planck 2013-2018 data, there has been a surge in the building of single-field Starobinsky-like inflationary scenarios [1] in various frameworks of quantum gravity due to stringent constraints on multifield inflation. The vanilla models of single-field inflation predict the scalar spectral index ($n_{s}$) and tensor-to-scalar ratio ($r$) as [2, 3]
@@ I Introduction
where $\alpha$ is a parameter that depends on the framework for the ultra-violet (UV) completion of gravity. In the context of $\alpha$-attractor inflation, the $\alpha$ parameter indicates the curvature of the Kähler geometry in supergravity [4, 5, 6, 7, 8, 9, 10, 11, 12]. It is important to note that models of $\alpha$-attractors could predict, in general, any smaller value for the tensor-to-scalar ratio compatible with the latest bound $r<0.032$ [13, 14]. Although future observations aim to detect primordial gravitational waves, there is a lack of a theoretical lower bound on the value of $\alpha$ except through considerations of reheating [10]. However, right after inflation ends, the universe enters a phase known as preheating [15, 16, 17, 18, 19] during which the universe behaves effectively close to matter-dominated phase with instabilities [20, 21, 22, 23]. Furthermore, when the inflationary potential deviates significantly from quadratic behavior (as is the case for small $\alpha$), the oscillations trigger the phenomenon of self-resonance (either by including or excluding metric fluctuations) [22, 23, 24, 25, 26, 20, 27, 28, 21, 29, 30, 31, 32, 33, 34, 35], which translates into a fast and substantial amplification of scalar perturbations that act as a source of Gravitational Waves (GWs), since these fluctuations couple to tensor ones at second order in the equations of motion or 3rd order in the perturbation of the action. The GWs produced in this fashion are called Scalar-Induced Gravitational Waves (SIGWs) [36, 37, 38, 39, 40, 41, 42, 43], and can be so efficiently produced to reach the Big Bang Nucleosynthesis (BBN) bound. This precise fact is what our investigation utilizes to derive a new lower bound on $\alpha$, based solely on the considerations of preheating instabilities and the consequent generation of SIGWs. Our result is robust even for a small duration of preheating of 1-2 e-folds, as one can realistically expect after the end of inflation.
@@
Recently, we have shown that these instabilities during preheating could lead to the generation of primordial black holes (PBH) [22, 44] in Starobinsky as well as in $\alpha$-attractor inflationary models. However, the study in this paper focuses on SIGWs produced irrespective of whether or not PBH are generated during preheating. It is also important to note that SIGWs are a secondary effect. At the linearized level, both scalar and tensor perturbations decouple. Therefore, we compute the linearized effect, which we call background GWs (BGWs), and the non-linear (secondary) effect, i.e., the SIGWs.
@@
Figure 1: Schematic representation of the production of SIGWs during self-resonant preheating. The upper panel shows the evolution of selected comoving scales as a function of the scale factor of the universe, as well as the comoving Hubble radius $(aH)^{-1}$. The suffixes CMB, end, and rh correspond to the CMB pivot scale, the end of inflation, and the beginning of reheating, respectively. The evolution of the Fourier mode of the metric fluctuation $\Phi_{k}$ in the bottom panel is shown schematically to illustrate the approximate behavior of the perturbations. See the text for details.
@@
A schematic representation of the scenario described above is shown in Fig. 1. The key aspect is the definition of the relevant scales involved in the instability, labeled as band of instability. This band is bounded by two characteristic scales: $k_{L}$ (lower) and $k_{U}$ (upper). The lower bound, $k_{L}$, corresponds to the last mode entering the horizon during preheating and is defined as $k_{L}=a(t_{\text{rh}})H(t_{\text{rh}})$, where $a(t)$ is the scale factor of the universe, $H(t)$ the Hubble rate, and $t_{\text{rh}}$ marks the end of the preheating phase. The upper bound, $k_{U}$, denotes the largest mode affected by self-resonance and is determined using Floquet theory [23, 21]. When the modes enter the self-resonance instability (green shaded area), the scalar perturbations, such as $\Phi_{\bm{k}}$, amplify. Their behavior is illustrated in the bottom panel of Fig. 1 for two distinct scales, $k_{1}$ and $k_{2}$, separated by the scale $k_{\text{end}}=a_{\text{end}}H_{\text{end}}$, defined as the last scale to exit the horizon during inflation (here and in what follows, the suffix "end" refers to evaluation at the end of inflation). The first scale, shown in red, remains approximately constant after re-entering the Hubble horizon $(aH)^{-1}$ during preheating. Although this may appear as non-amplification, it actually reflects the counteraction of the usual decay of the gravitational potential at sub-horizon scales in an expanding universe, ensuring $\Phi_{\bm{k}}$ remains stable during this phase. The mode $k_{2}$, shown in blue, experiences stronger amplification as it spends more time within the region of instability of self-resonance. It should be noted that the Floquet exponent, responsible for amplification, is not the same for all modes within the band of instability [23]. In fact, the amplification is particularly stronger for modes $k>k_{\text{end}}$. During this amplification process, SIGWs are generated, and after that, the SIGWs freely propagate throughout the universe with minimal interaction with matter.
@@
Depending on the specific range of scales affected by the self-resonance, the SIGWs will have an imprint in a particular span of frequencies. In general, these scales are small and around the last scale to exit the horizon during inflation, which implies that we lie on the Very-High-Frequency (VHF) region of the GW spectrum (100 kHz – 1 THz) [45, 46, 47, 48]. This VHF region has been explored in the literature in the contexts of GUT scale PBH and quantum gravity [49, 50, 51]. Although some detectors (in operation and/or planned) work on the VHF band of the GW spectrum, their sensitivities are too low to detect a stochastic background of SIGWs. These sensitivities are shown in Fig. 2 in terms of the fractional energy density of gravitational waves, $\Omega_{\text{GW}}$. One can observe that all of them lie above the cosmological bound coming from Big Bang nucleosynthesis (BBN) [52, 53], making it impossible to detect a stochastic background of GWs formed during preheating (otherwise, it would conflict with BBN predictions). In the VHF band, one of interest in this work, the most promising probes to detect GWs are:
@@
Figure 2: Sensitivity curves of some GW detectors (in terms of $\Omega_{GW}$) as a function of the frequency in Hertz. Also shown is the BBN bound, taken from [52, 53]. Solid (dashed) lines represent operative (planned or theorized) GW detectors. The figure is produced from the analysis of [45, 46] and references therein. The bottom lines with arrows mark the limits of the standard frequency bands of GWs, as defined in [45]. The blue and magenta curves labeled as $r=0.028$ and $r=2.14\times 10^{-6}$ represent two spectra of GWs computed from the upper and lower bounds on the tensor-to-scalar ratio; see the text for details. These are computed using the method outlined in this work.
  • Optically levitated sensors (OLS): These are based on the use of optically trapped and cooled dielectric microspheres or microdisks that experience a force when a GW passes through [54].
  • Bulk acoustic wave (BAW) devices: The vibrations of a resonant mass detector due to the passage of a GW could be read through specific devices [55, 56].
  • Interferometers up to 100 MHz: Laser interferometers measure differential arm-length changes induced by GWs. The quantum Cramér-Rao bound implies that, for a given laser power, higher bandwidth is needed to increase the sensitivity [57]. This implies that a broadband interferometer with the Virgo, LIGO, or KAGRA-level sensitivity is not viable in the MHz region. However, there are several other interferometers in this bandwidth, such as the Holometer experiment at Fermilab [58].
  • Gertsenshtein effect: This consists of the conversion of photons into GWs (and vice versa) in the presence of a magnetic field. Although these kinds of detectors are not yet built, Ref. [46] points out the possibility of using existing experiments, such as IAXO-SPD and IAXO-HET [59]. Another possibility is to rely on observations of large-scale cosmic regions with radio telescopes, such as EDGES or ARCADE [60, 61].
Apart from the VHF band, we also show for comparison in Fig. 2 the sensitivity curves of GW detectors in lower frequency bands: the Very-Low-Frequency (VLF), the Low-Frequency (LF), the Middle-Frequency (MF), and the High-Frequency (HF). See [45, 54] and references therein for details about the detectors and the acronyms. The two curves labeled as $r=0.028$ and $r=2.25\times 10^{-6}$ in Fig. 2 correspond to the energy density of GWs (background ++ induced) from an E-model with $\alpha=8.4$ and $\alpha=10^{-3.17}$, respectively, where $r$ is the tensor to scalar ratio (1). The upper bound on $r$ corresponds to the one obtained using 10 datasets from the BICEP/Keck Array 2015 and 2018, Planck releases 3 and 4, and LIGO-Virgo-KAGRA Collaboration [62]. The lower bound is obtained in this work using SIGWs and the BBN bound. See Sec. III.2 for details.
@@
The inflationary models we study in this work are one of the well-known universality-class of inflationary potentials [7], such as the $\alpha$-attractor models [4, 5, 6, 8, 9, 10, 11, 12], which exhibit excellent agreement with Planck data [63] and have universal predictions of observables. For this reason, they have gained a lot of attention. There are two main subclasses of $\alpha$-attractor models, called T-models and E-models, named by the shape of the inflaton potentials (see Fig. 3), which are given by
@@
$\begin{split}V_{T}(\phi)&=3\alpha M^{2}M_{\mathrm{Pl}}^{2}\tanh^{2}\left[\frac{1}{\sqrt{6\alpha}}\phi\right],\\ V_{E}(\phi)&=\frac{3\alpha M^{2}M_{\mathrm{Pl}}^{2}}{4}\left(1-e^{-\sqrt{\frac{2}{3\alpha}}\phi}\right)^{2},\end{split}$ (2)
@@
where $M_{\mathrm{Pl}}$ is the Planck mass and the scalar field mass $M=1.3\times 10^{-5}M_{\mathrm{Pl}}$ is fixed by the amplitude of primordial power spectra. Beyond their inflationary observables (1), we show that these models are also particularly interesting from the perspective of SIGWs. As the value of $\alpha$ decreases, the minimum of the potential (at $\phi=0$) starts to deviate from the quadratic (parabola) behavior, triggering the self-resonance phenomenon and amplifying the scalar perturbations. Here we restrict our study to (2) potentials. Fig. 3 compares both potentials in Eqn. (2) for several values of $\alpha$, revealing an evident asymmetry in the E-model case. This, as shown in [23], makes the self-resonance effect stronger in this class of potentials, ultimately imposing a tighter constraint on $\alpha$ for the E-models and, consequently, on the tensor-to-scalar ratio. These differences will be further analyzed in Sec. III, where their impact on tensor perturbations is clarified.
It would be non-trivial to extend our results to higher order $T,E$- shaped potentials of universality class [7]. For example, in the case of a potential with a quartic minimum (that comes from potential class $V_{T}\sim\tanh^{4}\left(\frac{\phi}{\sqrt{6\alpha}}\right)$ [7]), the perturbations do amplify, as shown in [18, 64, 65, 26]. However, as a consequence of this, the resonance band is likely to get narrowed compared to the potentials in (2). Although we do not show the results for this class of potentials, we expect this to affect the production of SIGWs in two ways. First, the universe in this case behaves effectively as radiation-dominated right after inflation, affecting the production of SIGWs, and second, this production, if any, would correspond to a narrower range of frequencies in comparison, but could also result in a stronger spectrum of GWs depending on the details of inflaton dynamics prior to its decay. The quantification of this effect would require further investigation, which goes beyond the scope of the paper.
@@
This work is organized as follows: In Sec. II, we analyze the evolution of scalar metric perturbations in a perturbed flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric. In Sec. III, we examine the generation of SIGWs from these scalar perturbations and numerically compute the energy density of GWs for different values of $\alpha$. Finally, conclusions are drawn in Sec . IV, and Appendixes A and B show some complementary computations, such as the derivation of the equation of motion for the SIGWs or its power spectrum. Throughout the paper, we follow the metric signature (+++)(-+++) and $\hbar=c=1$ with reduced Planck mass $M_{\rm Pl}^{2}=\frac{1}{8\pi G}$. Throughout the paper, overdots denote differentiation with respect to cosmic time ($t$) and overprimes denote differentiation with respect to conformal time ($\eta$).
(a) T-model
(b) E-model
@@
Figure 3: Normalized $\alpha$-attractor potentials for several values of $\alpha$: $10^{-1}$ (blue), $10^{-2}$ (orange), $10^{-3}$ (green), and $10^{-4}$ (red).
II Scalar perturbations